40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Sep 18 01:14:23 2008, rob, Summary, Computers, EPICS BAD 
    Reply  Thu Sep 18 12:16:01 2008, Yoichi, Update, Computers, EPICS BACK 
       Reply  Thu Sep 18 13:05:05 2008, Yoichi, Update, Computers, EPICS BACK 
Message ID: 961     Entry time: Thu Sep 18 01:14:23 2008     Reply to this: 963
Author: rob 
Type: Summary 
Category: Computers 
Subject: EPICS BAD 

Somehow the EPICS system got hosed tonight. We're pretty much dead in the water till we can get it sorted.

The alignment scripts were not working: the SUS_[opt]_[dof]_COMM CA clients were having consistent network failures.
I figured it might be related to the network work going on recently--I tried rebooting the c1susaux (the EPICS VME
processor in 1Y5 which controls all the vertex angle biases and watchdogs). This machine didn't come back after
multiple attempts at keying the crate and pressing the reset button. All the other cards in the crate are displaying
red FAIL lights. The MEDM screens which show channels from this processor are white. It appears that the default
watchdog switch position is OFF, so the suspensions are not receiving any control signals. I've left the damping loops
off for now. I'm not sure what's going on, as there's no way to plug in a monitor and see why the processor is not coming up.

A bit later, the c1psl also stopped communicating with MEDM, so all the screens with PSL controls are also white. I didn't try
rebooting that one, so all the switches are still in their nominal state.
ELOG V3.1.3-